Deborah, Dave and Judy are practicing their piano lessons. Deb practices as long as Dave and Judy practices twice as long as Dave. If their combined practice time is 110 minutes, how long does each practice.
Answer provided:
Dave 30 minutes, Deb 20 minutes, Judy 60 minutes
If Deb practices as long as Dave wouldn't their minutes be the same?
I put an x for Dave and 2x for Judy but don't know what to put for Deborah?
Can you help with this?
Thanks in advance.

Answer by jorel1380(3719)

Let Deborah and Dave be n. Then Judy would be 2n, since Judy practices twice as long as Deborah or Dave. Then:
n+n+2n=110
4n=110
n=110/4
2n=110/2
To get the results from the provided answer, Deb would have to practice 2/3 of what Dave does
☺☺☺☺